There are 6 playing cards placed face down. the 4 of diamond , the 5 of hearts, the 2 of clubs, the 10 of spades, the 7 of hearts and the 3 diamonds. A person selects a card , find the expected value of ther dra.
Answer by stanbon(75887)

There are 6 playing cards placed face down.
the 4 of diamond ,
the 5 of hearts,
the 2 of clubs,
the 10 of spades,
the 7 of hearts and
the 3 diamonds.
A person selects a card , find the expected value of their draw.
-----
The probability of drawing each card is (1/6)
--
Multiply the "value" of each card by its probability.
--
Add the 6 products to get the expected value.
---------
E(x) = (1/6)(4+5+2+10+7+3) = 31/6 = 5.167
==================
Cheers,
Stan H.
==================